OpenStudy (destinyyyy):

Can someone explain how to solve this? Thank you. Find the coordinates of the vertex and the intercepts of the following quadratic function. When necessary, approximate the x-intercepts to the nearest tenth.

OpenStudy (destinyyyy):

f(x)+x^2+6x+5

OpenStudy (destinyyyy):

a= 1 b= 6 c= 5 x= -(6)/2(1) =-6/2 =-3

OpenStudy (destinyyyy):

@pooja195

OpenStudy (princeharryyy):

that's the x coordinate of the vertex.

OpenStudy (princeharryyy):

to find the y coordinate you need to put x =-3 in the equation and solve for the value of y.

OpenStudy (destinyyyy):

f(-3)=(-3)^2 -6(-3)+5 = 9+18+5 32

OpenStudy (princeharryyy):

where y is nothing but f(x) itself. and find the ordered pair as V(x,f(x))

OpenStudy (princeharryyy):

f(x)=x^2+6x+5 f(-3) = (-3)^2 +6*(-3) +5 = 9-18+5 =4

OpenStudy (destinyyyy):

vertex= (-3,-4) ?

OpenStudy (princeharryyy):

yeah that will be -4 not 4 great, yeah! that's your vertex.

OpenStudy (destinyyyy):

Okay... For y-intercept f(0)=(0)^2+6(0)+5 5 (0,5) ?

OpenStudy (princeharryyy):

yes!

OpenStudy (destinyyyy):

Okay

OpenStudy (princeharryyy):

for x intercepts you need to put x^2+6x+5 = 0 and find all possible x's.

OpenStudy (destinyyyy):

Okay so I have to factor x^2+6x+5=0

OpenStudy (princeharryyy):

yes! (x+1)(x+5) =0; x=-1,-5 and f(x)'s at these x's are 0.

OpenStudy (destinyyyy):

Okay.. Is that everything?

OpenStudy (princeharryyy):

For this question.

OpenStudy (destinyyyy):

Okay... Thank you.

OpenStudy (princeharryyy):

:) wlcm